Identify the following conic and hence reduce it to standard form:
36x² + 20xy + 30y² − 70x + 128y + 80 = 0
1
Expert's answer
2020-09-17T13:38:01-0400

The expression for a conic section in the Cartesian coordinate system is defined as:


"Ax^2+Bxy+Cy^2+Dx+Ey+F=0, A\not=0, B\not=0,C\not=0"

The result of "B^2-4AC"  determines the type of the conic section obtained.



"36x^2+20xy+30y^2-70x+128y+80=0""18x^2+10xy+15y^2-35x+64y+40=0"

"A=18, B=10, C=15, D=-35, E=64,F=40"

"B^2-4AC=(10)^2-4(18)(15)=-980<0"

Then the curve is ellipse, circle, point or no curve.

A. Rotate to remove xy-term

Rotate the coordinate axes counterclockwise through the angle "\theta" where "\theta" is defined by "\cot(2\theta)=\dfrac{A-C}{B}=\dfrac{18-15}{10}=0.3=\dfrac{\cot^2\theta-1}{\cot\theta}"



"\cot^2\theta-0.3\cot\theta-1=0""\cot\theta=\dfrac{0.3\pm\sqrt{0.09+4}}{2}=\dfrac{0.3\pm\sqrt{4.09}}{2}"

Take "\cot\theta=\dfrac{0.3+\sqrt{4.09}}{2}"

"x=x'\cos \theta-y'\sin \theta"

"y=x'\sin \theta+y'\cos \theta"



"18(x')^2\cos^2 \theta-36x'y'\sin\theta\cos\theta+18(y')^2\sin^2\theta+""+10(x')^2\sin\theta\cos\theta+10x'y'\cos^2\theta-10x'y'\sin^2\theta-""-10(y')^2\sin\theta\cos\theta+15(x')^2\sin^2\theta+30x'y'\sin\theta\cos\theta+""+15(y')^2\cos^2\theta-35x'\cos\theta+35y'\sin\theta+""+64x'\sin \theta+64y'\cos \theta+40=0""(x')^2(18\cos^2 \theta+10\sin\theta\cos\theta+15\sin^2\theta)+""+(y')^2(18\sin^2 \theta-10\sin\theta\cos\theta+15\cos^2\theta)+""+x'(64\sin\theta-35\cos\theta)+y'(64\cos\theta+35\sin\theta)+40=0""-36\sin\theta\cos\theta+10\cos^2\theta-10\sin^2\theta+30\sin\theta\cos\theta=""=10\cos(2\theta)-3\sin(2\theta)=10\cos(2\theta)(1-3\cot(2\theta))=0""(15+3\cos^2 \theta+10\sin\theta\cos\theta)(x')^2+""+(15+3\sin^2 \theta-10\sin\theta\cos\theta)(y')^2+""+64x'\sin \theta+64y'\cos \theta+40=0""(15+3\cos^2 \theta+5\sin(2\theta))(x'+\dfrac{32\sin\theta}{15+3\cos^2 \theta+5\sin(2\theta)})^2+""+(15+3\sin^2 \theta-5\sin(2\theta))(y'+\dfrac{32\cos\theta}{15+3\sin^2 \theta-5\sin(2\theta)})^2=""=\dfrac{1024\sin^2\theta}{15+3\cos^2 \theta+5\sin(2\theta)}+""+\dfrac{1024\cos^2\theta}{15+3\sin^2 \theta-5\sin(2\theta)}-40"

"u=\dfrac{32\sin\theta}{15+3\cos^2 \theta+5\sin(2\theta)}"

"v=\dfrac{32\cos\theta}{15+3\sin^2 \theta-5\sin(2\theta)}"

"c^2=\dfrac{1024\sin^2\theta}{15+3\cos^2 \theta+5\sin(2\theta)}+"

"+\dfrac{1024\cos^2\theta}{15+3\sin^2 \theta-5\sin(2\theta)}-40"


"a^2=(15+3\cos^2 \theta+5\sin(2\theta))"

"\dfrac{15360+3072\sin^4\theta+3072\cos^4\theta+2560\sin(2\theta)}{(15+3\cos^2 \theta+5\sin(2\theta))(15+3\sin^2 \theta-5\sin(2\theta))}"

"b^2=15+3\sin^2 \theta-5\sin(2\theta)"





"\dfrac{(x'+u)^2}{\dfrac{c^2}{a^2}}+\dfrac{(y'+v)^2}{\dfrac{c^2}{b^2}}=1"

Ellipse.